# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ 3 Bias Correction and Asymptotic Theory
$[\widehat{H}(\theta)]^{-1}=:\left(\begin{array}[c]{cc}\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ \widehat{H}_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right)$ (3.3)
in the same manner as in (3.2). The term $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ can be estimated by the $(N-1)K\times(N-1)K$ block $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, and $H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$ by the $(T-1)K\times(T-1)K$ block $\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$. Estimators of $S_{\alpha\alpha}(\theta)$ and $S_{\gamma\gamma}(\theta)$ are constructed from the derivatives of $l_{i,t}(\theta,\phi)$, the original likelihood for individual $i$ at time $t$. This procedure needs some explanation. The reparameterization ensures the following identity:
where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)=D^{\prime}\psi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t. $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$. Specifically,
$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)),\qquad s_{i,t}^{\alpha}(\theta,\phi):=\partial_{\alpha_{i}}l_{i,t}(\theta,\phi).$
Here, $\widetilde{s}_{i,t}^{\alpha}(\theta)$ is the centered version of $s_{i,t}^{\alpha}(\theta,\phi(\theta))$, the derivative, w.r.t. $\alpha_{i}$ and evaluated at $\phi(\theta)$, of the original likelihood $l_{i,t}(\theta,\phi)$ for individual $i$ at time $t$. Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
Naturally, the estimator of $S_{\alpha\alpha}(\theta)$ is
$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta)).$
The indicator $\mathds{1}\{\left|t-s\right|\leq\tau\}$ is a truncation mechanism with truncation parameter $\tau$ (see Remark 3 below). Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)),\qquad s_{i,t}^{\gamma}(\theta,\phi):=\partial_{\gamma_{t}}l_{i,t}(\theta,\phi),$
where $\widetilde{s}_{i,t}^{\gamma}(\theta)$ is the centered version of $s_{i,t}^{\gamma}(\theta,\phi(\theta))$, the derivative of $l_{i,t}(\theta,\phi)$ w.r.t. $\gamma_{t}$ and evaluated at $\phi(\theta)$. By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (3.5)
Intuitively, the estimator for $S_{\gamma\gamma}(\theta)$ is
$\widehat{S}_{\gamma\gamma}(\theta)$ $=D_{2}\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)D_{2}^{\prime},$ (3.6)
$\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}\widehat{s}_{i}^{\gamma}(\theta)[\widehat{s}_{i}^{\gamma}(\theta)]^{\prime}\},\qquad\widehat{s}_{i}^{\gamma}(\theta):=\{[\widehat{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widehat{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta)).$
Remark 3 (truncation)
In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016). This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case. Detailed discussions in this regard can be found in Andrews (1991). Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2). Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2 below). Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$. For intermediate $T$, we recommend following the advices from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values. In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (3.5) using cross-sectional independence. One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 00. We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics. The comparison is, hence, not reported in the paper.
Using Equations (3.3), (3.4), and (3.6), we can construct the corrected likelihood $L(\theta)$, which is stated in the theorem below with its rate of convergence.
Theorem 2
Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta),$
as in (3.1), with
As mentioned above, Theorem 2 states that $L(\theta)$ does not suffer from the IPP. Showing 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1. Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error. For the truncation, we show that $1/\left(N^{2}T^{2}\right)\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\cdot\right)$, which is the truncated (i.e., the removed) part in (3.4), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions. Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
@@
Next, we turn our attention to the asymptotic properties of the $L(\theta)$-based estimators $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta)$ and of the $L(\theta)$-based LR, LM, Wald test statistics. We enforce the following assumption.
@@
Here Assumption 2 is similar to Assumption 1, except that we impose that $l_{i,t}(\theta,\phi)$ is four-time differentiable, and require that the derivatives of $l_{i,t}(\theta,\phi)$ w.r.t. to both $\theta$ and $\phi$ to be uniformly dominated by $g(w_{it})$. This assumption ensures that the corrected likelihood $L(\theta)$, which depends on second derivatives of $l_{i,t}(\theta,\phi)$, is twice differentiable. This is required by standard techniques used to show, e.g., the asymptotic normality of $\widehat{\theta}_{L}$. See., e.g., Hayashi (2000). Similar to Assumption 1, Assumption 2 is an identification condition for $\theta_{0}$. It guarantees that $\operatorname*{plim}_{N,T\rightarrow\infty}l(\theta)$ is uniquely maximized at $\theta_{0}$. Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 00 uniformly as $N,T\rightarrow\infty$. Together, Assumptions 2 to 2 are used to establish the consistency of $\widehat{\theta}_{L}$. In addition to these, Assumption 2 establishes a dominance condition for the profiled Hessian, so that $\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})\overset{\mathbb{P}}{\longrightarrow}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ for any consistent estimator $\overline{\theta}$ of $\theta_{0}$. Assumption 2 requires that the expected profiled Hessian is strictly negative definite for sufficiently large $N$ and $T$. Assumptions 2 and 2 are additionally imposed to establish the asymptotic distributions for the quantities of interests.
@@
Here, $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$. By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues. In addition, since we focus on maximum likelihood models, we use the information matrix equality to simplify the presentation of Corollary 1. In case the information matrix equality does not hold [^6] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof. See, e.g., White (1982) for detailed discussions.
For inference procedures, we consider a generic null hypothesis $H_{0}:R(\theta_{0})=0$, where $R(\theta)$ is a known $r\times 1$ ($r\leq K$) vector-valued non-random function with Jacobian $J(\theta):=\partial_{\theta^{\prime}}R(\theta)$ satisfying $\operatorname*{rank}[J(\theta)]=r$. The corollary below gives the asymptotic distribution, under $H_{0}$, of $\widehat{\xi}_{LR}$ (the LR statistic), $\widehat{\xi}_{LM}$ (the LM statistic), and $\widehat{\xi}_{Wald}$ (the Wald statistic), defined respectively as
where $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0$.
@@
as $N,T\rightarrow\infty$, where $\chi^{2}(r)$ is the $\chi^{2}$-distribution with degrees of freedom $r$.
Similarly, the test statistics $\widehat{\xi}_{LM}$ and $\widehat{\xi}_{Wald}$ are constructed under the information matrix equality. Adjustments are generally required if it does not hold. For instance, White (1982) shows that $-[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.

### v5

@@ C.2 Regularity Conditions
$X^{\gamma}=\left(\begin{array}[c]{ccc}X_{11}^{\prime}&&\\ &\ddots&\\ &&X_{1T}^{\prime}\\ X_{21}^{\prime}&&\\ &\ddots&\\ &&X_{2T}^{\prime}\\ \vdots&\ddots&\vdots\\ X_{N1}^{\prime}&&\\ &\ddots&\\ &&X_{NT}^{\prime}\end{array}\right),\qquad X^{\alpha}=\left(\begin{array}[c]{cccc}X_{11}^{\prime}&&&\\ \vdots&&&\\ X_{1T}^{\prime}&&&\\ &X_{21}^{\prime}&&\\ &\vdots&&\\ &X_{2T}^{\prime}&&\\ &&\ddots&\\ &&&X_{N1}^{\prime}\\ &&&\\ &&&X_{NT}^{\prime}\end{array}\right).$
Moreover, $H^{\pi}(\theta)=\operatorname*{diag}\left\{H_{1}^{\pi}(\theta),\ldots,H_{N}^{\pi}(\theta)\right\}$ is a diagnal matrix, where block $H_{i}^{\pi}(\theta)=\operatorname*{diag}\left\{h_{i,1}^{\pi}(\theta),\ldots,h_{i,T}^{\pi}(\theta)\right\}$ and each element is given by
$h_{i,t}^{\pi}(\theta)=-\frac{\exp(-\pi_{i,t}(\theta))}{[1+\exp(-\pi_{i,t}(\theta))]^{2}}<0,\quad\pi_{i,t}(\theta)=X_{it}^{\prime}(\theta+\alpha_{i}(\theta)+\gamma_{t}(\theta)).$
Here, $\alpha_{i}(\theta)$ and $\gamma_{t}(\theta)$ denote the pseudo-true values defined by (2.4). Let $\lambda_{\max}(A)$ denote the maximum eigenvalue of a matrix $A.$ Using (C.6), we have
$\lambda_{\max}(\sqrt{NT}\mathbb{E}H(\theta))\leq\sup_{i,t}h_{i,t}^{\pi}(\theta)\lambda_{\max}(DX^{\prime}XD^{\prime}/\sqrt{NT}).$
Therefore, Assumption 1 is implied if $XD^{\prime}$ has full column rank. Given the structure of $X$, this requires that, for all $i$, the columns of $X_{i\cdot}$ are linearly independent and that, for all $t$, the columns of $X_{\cdot t}$ are linearly independent.
The first condition requires that certain higher moment of $\left\|X_{it}\right\|_{\max}$ are bounded away from the infinity uniformly. Note that $X_{it}$ has a fixed dimension, so that the choice of norm is irrelevant. This is relatively a mild condition. The second condition states that the regressors are linearly independent in both the individual and time dimension. This condition is standard for models with two-way heterogeneities, and it is usually satisfied if the regressors have sufficient variations across both the cross-sectional and time dimension. See Remark 1 of Fernández-Val and Weidner (2016) for a similar argument. Notably, it rules out the possibility of $Y_{it-1}$ being constant across $i$ or $t$.
@@
To establish the asymptotic properties of $\widehat{\theta}_{L}$ and $L(\theta)$-based LR, LM, Wald test statistics, we enforce the following assumption.
@@
    where $\left\|\cdot\right\|$ is the matrix Frobenius norm.
Here Assumption 2 is similar to Assumption 1, except that we impose that $l_{i,t}(\theta,\phi)$ is four-time differentiable, and require that the derivatives of $l_{i,t}(\theta,\phi)$ w.r.t. to both $\theta$ and $\phi$ to be uniformly dominated by $g(w_{it})$. This assumption ensures that the corrected likelihood $L(\theta)$, which depends on second derivatives of $l_{i,t}(\theta,\phi)$, is twice differentiable. This is required by standard techniques used to show, e.g., the asymptotic normality of $\widehat{\theta}_{L}$. See., e.g., Hayashi (2000). Similar to Assumption 1, Assumption 2 is an identification condition for $\theta_{0}$. It guarantees that $\operatorname*{plim}_{N,T\rightarrow\infty}l(\theta)$ is uniquely maximized at $\theta_{0}$. Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 0 uniformly as $N,T\rightarrow\infty$. Together, Assumptions 2 to 2 are used to establish the consistency of $\widehat{\theta}_{L}$. In addition to these, Assumption 2 establishes a dominance condition for the profiled Hessian, so that $\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})\overset{\mathbb{P}}{\longrightarrow}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ for any consistent estimator $\overline{\theta}$ of $\theta_{0}$. Assumption 2 requires that the expected profiled Hessian is strictly negative definite for sufficiently large $N$ and $T$. Assumptions 2 and 2 are additionally imposed to establish the asymptotic distributions for the quantities of interests.
C.3 Asymptotic Theory
Theorem 1
Under Assumption 1 and uniformly over $\theta\in\Theta$,
as $N,T\rightarrow\infty$, where
$B_{\alpha}(\theta)=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta)=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
Proof. In Section E.
@@ C.3 Asymptotic Theory
The next theorem, which shows that approximation error of the corrected likelihood $L(\theta)$ to the infeasible IPP-free likelihood $l(\theta)$ is of order $1/\sqrt{NT}$ uniformly over $\theta$. This is our claim in the beginning of this section.
Theorem 2
Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$, we have
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where $L(\theta)$ is the corrected likelihood defined in Equation (3.3).
Showing Theorem 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1. Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error. For the truncation, we show that
$\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\right),$
which is the truncated (i.e., the removed) part in (3.6), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions. Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
@@
Proof. In Section E.
Here, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$. By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues. In addition, since we focus on maximum likelihood models, we use the information matrix equality to simplify the presentation of Corollary 1. In case the information matrix equality does not hold [^7] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof. See, e.g., White (1982) for detailed discussions.
The next corollary states the asymptotic distribution of the LR ($\widehat{\xi}_{LR}$), the LM ($\widehat{\xi}_{LM}$), and the Wald ($\widehat{\xi}_{Wald}$) statistics.
@@
as $N,T\rightarrow\infty$.
Proof. In Section E.
Similarly, the test statistics $\widehat{\xi}_{LM}$ and $\widehat{\xi}_{Wald}$ are constructed under the information matrix equality. Adjustments are generally required if it does not hold. For instance, White (1982) shows that $\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.